OpenStudy (anonymous):

Find the approximate solution of the given system of equations. 8x + 9y = 9 3x – 5y = 20 A. (3.36, –1.99) B. (–1.11, 2.79) C. (0, –1.99) D. (3.36, 2.79)

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

it A...

OpenStudy (texaschic101):

8x + 9y = 9 -->(-3) 3x - 5y = 20 -->(8) ----------------- -24x - 27y = - 27 (result of multiplying by -3) 24x - 40y = 160 (result of multiplying by 8) -----------------add - 67y = 133 y = -133/67 y = - 1.985 ....rounds to -1.99 now just sub -1.99 in for y in either equation to find x zaibali is correct...it is A